OpenStudy (anonymous):

Tree enthusiasts use a point system to compare trees. They assign a number of points for each inch of circumference, foot of height, and foot of crown spread. The table shows the points in each category and the total points for three of South Carolina's trees. How many points are assigned to one unit of each measurement? (Hint: Write a system of three equations and solve for each variable.) The Table: Circumference(in): Height(ft): Crown Spread (ft): Total Points Angel Oak:306: 65: 144: 407 Green Ash: 181: 143: 96: 348 Redwood: 168: 105: 44: 284

OpenStudy (anonymous):

Let points assigned to each inch of circumference be x, to each foot of height be y and to each foot of crown spread be x. Then, 306x + 65y + 144z = 407 similarly, 181x + 143y + 96 z = 348 168x + 105y +44z = 248 You have 3 variables 3 equations... solve them to get x,y and z.

OpenStudy (anonymous):

a. circumference: 1 point; height: 1 point; crown spread: 0.75 point b. circumference: 2 point;s height: 1 point; crown spread: 0.25 point c. circumference: 1 point; height: 2 points; crown spread: 0.25 point d. circumference: 1 point; height: 1 point; crown spread: 0.25 point would it be c?